# Can Big Black Holes Merge with the Smallest Black Holes?

### v1

@@ 3.1 Evolution of minimum secondary mass
Under the Power Law minimum secondary-mass model, the evolution of $m_{2}^{\mathrm{min}}$ is preferentially flat until high primary masses, where the uncertainty on $m_{2}^{\mathrm{min}}$ grows. At high $m_{1}$, $m_{2}^{\mathrm{min}}$ is forced to increase to $m_{\rm{max}}$ because of the functional form of the model. The hyperposteriors in Figure 2 show that high $\gamma$ is preferred, meaning that the evolution of $m_{2}^{\mathrm{min}}$ is flatter with $m_{1}$, before sharply increasing as $m_{1}$ approaches $m_{\mathrm{max}}$. The results are restricted by the prior on $\gamma$, as we allow a maximum $\gamma$ of $35$, while a flat evolution in $m_{2}^{\mathrm{min}}$ would require $\gamma\rightarrow\infty$. With our prior choice, we find that $\gamma>4.4$ with GW190814 and $\gamma>2.8$ without GW190814 at $90\%$ credibility. The preference for a flatter evolution of $m_{2}^{\mathrm{min}}$ means that there is little support for a higher minimum secondary mass at higher primary masses. While this behaviour is consistent whether or not GW190814 is included, the overall minimum mass $m_{\mathrm{min}}$ and the primary mass power-law index $\alpha$ are different when including GW190814. Without GW190814, the minimum mass is higher, with $m_{\mathrm{min}}=5.8^{+0.2}_{-0.6}$, while with GW190814, $m_{\mathrm{min}}=2.2^{+0.4}_{-0.2}$ to accommodate GW190814’s secondary mass.
Figure 2: Hyperposteriors on the key mass and mass-ratio hyperparameters with Power Law model, with and without GW190814. The $m_{2}^{\mathrm{min}}$ hyperparameter, $\gamma$, prefers high values, supporting a non-increasing or shallowly increasing $m_{2}^{\mathrm{min}}$ with $m_{1}$. The other hyperparameters show disagreement between the choice to include or exclude GW190814.
@@
Figure 4: Hyperposteriors on the key mass and mass ratio hyperparameters with Relaxed Parabola model, with and without GW190814. The $m_{2}^{\mathrm{min}}$ hyperparameters, $\xi$ and $\zeta$ are consistent with zero without GW190814, but negative $\xi$ is preferred when including GW190814. The other hyperparameters are in agreement between the two choices of observations.
We compare each of our analyses to the results with the default Power Law + Peak model from Abbott et al. (2023a), with $m_{2}^{\mathrm{min}}$ fixed to be equal to $m_{\mathrm{min}}$. We find that our models are generally not preferred to the default model without a variable $m_{2}^{\mathrm{min}}$, as shown in Table 2. This reinforces the result that there is no evidence that $m_{2}^{\mathrm{min}}$ (strictly) increases with $m_{1}$. The exception is the Relaxed Parabola model, which is preferred with a log Bayes factor of $\log_{10}\mathcal{B}=4.44$ when including GW190814. Therefore, if GW190814’s source is a BBH, this indicates that there is structure in the mass-ratio and component-mass distributions that should be accounted for by population models.
@@
Table 2: Log Bayes factors between the Power Law, Increasing Parabola, and Relaxed Parabola minimum secondary-mass models and the default model without evolution of the minimum secondary mass from Abbott et al. (2023a). All models allowing for an evolving minimum secondary mass with primary mass are disfavoured compared to the default model, with the exception of the Relaxed Parabola model when including GW190814.
@@ 3.2 Mass-ratio and primary-mass distributions
Figure 6: Slices of the median and $90\%$ interval for the mass-ratio distribution at primary masses of $10~M_{\odot}$ (left), $20~M_{\odot}$ (middle), and $40~M_{\odot}$ (right) for the Power Law (top), Increasing Parabola (middle), and Relaxed Parabola (bottom) models, each with and without GW190814. For the Increasing Parabola and Power Law models, the distributions are different when including GW190814, while for the Relaxed Parabola model the intervals are more consistent.
Compared to our analyses, Callister & Farr (2024b) and Sadiq et al. (2024) find that $p(q)$ has more support for unequal-mass mergers when considering data-driven models. These models do not consider a parametrised $m_{\mathrm{min}}$ and therefore have more flexibility to extend support to lower $q$ even at low $m_{1}$. Sadiq et al. (2024), when excluding GW190814, find that $p(q\mid m_{1}=10~M_{\odot})$ has support down to $q\sim 0.3$, while the distribution peaks away from equal masses at higher $m_{1}$.
@@ 4 Conclusions
We investigated the correlation between mass ratio and primary mass of merging BBHs to understand whether the biggest stellar-mass black holes can merge with the smallest. We introduced population models that included a minimum secondary mass $m_{2}^{\mathrm{min}}(m_{1})$ that varies with primary mass. We considered three model variations for $m_{2}^{\mathrm{min}}$, which we used with a power-law model for the mass-ratio distribution and a Power Law + Peak distribution for primary mass (Talbot & Thrane, 2018; Abbott et al., 2023a). Using these models we inferred how the minimum secondary mass evolves for GWTC-3.0 observations (Abbott et al., 2023b).
@@
While our models provide one way to understand correlations between primary mass and mass ratio, improvements to our model could be made to ensure robustness and avoid model misspecification. Our model that allows for decreasing $m_{2}^{\mathrm{min}}$ enforces that the minimum secondary mass is always positive, but currently does not place other bounds on $m_{2}^{\mathrm{min}}$, e.g., it could go to subsolar masses. Future work could update this assumption by imposing a non-zero lower bound on $m_{2}^{\mathrm{min}}$, representing a minimum black hole mass, which may help us understand the classification of GW190814. Additional model flexibility also could be added, for example by relaxing the assumption that $m_{2}^{\mathrm{min}}(m_{\mathrm{min}})=m_{\mathrm{min}}$; Farah et al. (2024) find that the inferred minimum primary and minimum secondary masses of the BBH population are in agreement within current uncertainties, but did not include GW190814 in their analysis. Our parametric forms for the minimum secondary mass are also not motivated by particular astrophysics or trend in the data. A data-driven description of the primary-mass and mass-ratio distributions may give more insight into the underlying correlations, unrestricted by the choice of parametric form. Furthermore, we specify a power-law distribution in mass ratio, restricting the mass-ratio distribution to be strictly monotonic. However, models with more relaxed assumptions about the shape of the distribution show support for the mass-ratio distribution to peak away from $1$ (Sadiq et al., 2024; Godfrey et al., 2023; Rinaldi et al., 2024; Callister & Farr, 2024a; Rinaldi et al., 2025). Considering a mass-ratio distribution that more accurately reflects the data would help avoid biases driven by an inaccurate models.
@@
Other observable parameters such as the spin or eccentricity could provide more information about the astrophysical nature of such BBHs. Works looking at the correlations between mass and spin may identify subpopulations of sources from similar evolutionary pathways (e.g., Safarzadeh et al., 2020; Heinzel et al., 2024; Ray et al., 2024; Antonini et al., 2025). If the mass-ratio distribution is jointly correlated with component mass and spin distributions, neglecting to model these correlations could lead to biased inference (Alvarez-Lopez et al., 2025). Jointly investigating mass, mass-ratio, and spin distributions could help us understand how these parameters are correlated, and therefore give more insight into the astrophysical evolution of the population. However, adding additional parameters to our models and investigating the joint correlations between mass, mass ratio, and spin, requires having sufficient data to constrain the model parameters. The increased number of BBH observations coming with the forth LIGO–Virgo–KAGRA observing run will allow us to better measure any correlations between the BBH observables, and explore the astrophysics encoded in these correlations.
@@ acknowledgments
We thank the groups at Northwestern University for hospitality while this work was developed, in particular to Sylvia Biscoveanu, Monica Gallegos-Garcia, and Miguel Martinez for insightful discussions. We thank Alexandra Guerrero for comprehensive review, Daniel Holz for valuable insights, and John Veitch for useful technical discussions. SC is supported by Science and Technology Facilities Council (STFC) studentship 2748218. CPLB is supported by STFC grant ST/V005634/1. ZD acknowledges support from the Center for Interdisciplinary Exploration and Research in Astrophysics (CIERA) Board of Visitors Research Professorship. SC thanks the CIERA Board of Visitors and the University of Glasgow MacRobertson Scholarship for support to visit CIERA. This research has made use of data from the Gravitational Wave Open Science Center, a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by the National Science Foundation (NSF) LIGO Laboratory which is a major facility fully funded by the NSF, as well as STFC of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO 600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by NSF Grants PHY-0757058 and PHY-0823459. This document has been assigned LIGO document number LIGO-P2500498.
These results were produced using models found in this code release. Results and data products used in this work are available on Zenodo: https://doi.org/10.5281/zenodo.16895784 (catalog doi:10.5281/zenodo.16895784) (Colloms et al., 2025b).

### v2

@@ 3.1 Evolution of minimum secondary mass
Under the Power Law minimum secondary-mass model, the evolution of $m_{2}^{\mathrm{min}}$ is preferentially flat until high primary masses, where the uncertainty on $m_{2}^{\mathrm{min}}$ grows. At high $m_{1}$, $m_{2}^{\mathrm{min}}$ is forced to increase to $m_{\mathrm{max}}$ because of the functional form of the model. The hyperposteriors in Figure 3 show that high $\gamma$ is preferred, meaning that the evolution of $m_{2}^{\mathrm{min}}$ is flatter with $m_{1}$, before sharply increasing as $m_{1}$ approaches $m_{\mathrm{max}}$. The results are restricted by the prior on $\gamma$, as we allow a maximum $\gamma$ of $35$, while a flat evolution in $m_{2}^{\mathrm{min}}$ would require $\gamma\rightarrow\infty$. With our prior choice, we find that $\gamma>4.4$ with GW190814 and $\gamma>2.8$ without GW190814 at $90\%$ credibility. The preference for a flatter evolution of $m_{2}^{\mathrm{min}}$ means that there is little support for a higher minimum secondary mass at higher primary masses. While this behaviour is consistent whether or not GW190814 is included, the overall minimum mass $m_{\mathrm{min}}$ and the primary-mass power-law index $\alpha$ are different when including GW190814. Without GW190814, the minimum mass is higher, with $m_{\mathrm{min}}=5.8^{+0.2}_{-0.6}~M_{\odot}$, while with GW190814, $m_{\mathrm{min}}=2.2^{+0.4}_{-0.2}~M_{\odot}$ to accommodate GW190814’s secondary mass.
Figure 3: Hyperposteriors on the key mass and mass-ratio hyperparameters with Power Law model, with and without GW190814. The $m_{2}^{\mathrm{min}}$ hyperparameter $\gamma$ prefers high values, supporting a non-increasing or shallowly increasing $m_{2}^{\mathrm{min}}$ with $m_{1}$. The other hyperparameters show disagreement between the choice to include or exclude GW190814.
@@
Figure 5: Hyperposteriors on the key mass and mass-ratio hyperparameters with the Relaxed Parabola model, with and without GW190814. The $m_{2}^{\mathrm{min}}$ hyperparameters, $\xi$ and $\zeta$ are consistent with zero without GW190814, but negative $\xi$ is preferred when including GW190814. The other hyperparameters are in agreement between the two choices of observations.
We compare each of our analyses to the results with the default Power Law + Peak model from KAGRA:2021duu, with $m_{2}^{\mathrm{min}}$ fixed to be equal to $m_{\mathrm{min}}$. We find that our models are generally not preferred to the default model without a variable $m_{2}^{\mathrm{min}}$, as shown in Table 2. This reinforces the result that there is no evidence that $m_{2}^{\mathrm{min}}$ (strictly) increases with $m_{1}$. The exception is the Relaxed Parabola model, which is preferred with a log Bayes factor of $\log_{10}\mathcal{B}=4.44$ when including GW190814. Therefore, if GW190814’s source is a BBH, this indicates that there is structure in the mass-ratio and component-mass distributions that should be accounted for by population models.
@@
Table 2: Log Bayes factors between the Power Law, Increasing Parabola, and Relaxed Parabola minimum secondary-mass models and the default model without evolution of the minimum secondary mass from KAGRA:2021duu. All models allowing for an evolving minimum secondary mass with primary mass are disfavoured compared to the default model, with the exception of the Relaxed Parabola model when including GW190814.
@@ 3.2 Mass-ratio and primary-mass distributions
Figure 7: Slices of the median and $90\%$ interval for the mass-ratio distribution at primary masses of $10~M_{\odot}$ (left), $20~M_{\odot}$ (middle), and $40~M_{\odot}$ (right) for the Power Law (top), Increasing Parabola (middle), and Relaxed Parabola (bottom) models, each with and without GW190814. The $99\%$ credible interval of the prior at each slice is shown in grey. For the Increasing Parabola and Power Law models, the distributions are different when including GW190814, while for the Relaxed Parabola model the intervals are more consistent.
Compared to our analyses, Callister:2023tgi and Sadiq:2023zee find that $p(q)$ has more support for unequal-mass mergers when considering data-driven models. These models do not consider a parametrised $m_{\mathrm{min}}$ and therefore have more flexibility to extend support to lower $q$ even at low $m_{1}$. Sadiq:2023zee, when excluding GW190814, find that $p(q\mid m_{1}=10~M_{\odot})$ has support down to $q\sim 0.3$, while the distribution peaks away from equal masses at higher $m_{1}$. Using parametric models, Li:2022jge find that higher-mass BBHs are more likely to be equal mass than lower-mass binaries when excluding GW190814. Meanwhile, Galaudage:2024meo, Li:2024jzi, and Roy:2025ktr find that BBHs around $m_{1}\sim 35~M_{\odot}$ are more likely to prefer equal masses. As shown in Figure 7, our models induce that the mass-ratio distribution varies with primary mass when excluding GW190814: at the lowest masses, the distribution looks steep because the range of potential mass ratios is tightly constrained by the minimum mass, while at larger masses the preferred distribution is comparatively flatter, though with broad uncertainties. The variation between results across the literature illustrates how different model constraints and prior assumptions can influence the interpretation of results, and highlights the importance of exploring alternative choices.
@@ 4 Conclusions
We investigated the correlation between the mass ratio and primary mass of merging BBHs to understand whether the biggest stellar-mass black holes can merge with the smallest. We introduced population models that included a minimum secondary mass $m_{2}^{\mathrm{min}}(m_{1})$ that varies with primary mass. We considered three model variations for $m_{2}^{\mathrm{min}}$, which we used with a power-law model for the mass-ratio distribution and a Power Law + Peak distribution for primary mass (Talbot:2018cva; KAGRA:2021duu). Using these models we inferred how the minimum secondary mass evolves for GWTC-3.0 observations (LIGOScientific:2021djp).
@@
While our models provide one way to understand correlations between primary mass and mass ratio, improvements to our model could be made to ensure robustness and avoid model misspecification. Our model that allows for decreasing $m_{2}^{\mathrm{min}}$ enforces that the minimum secondary mass is always positive, but currently does not place other bounds on $m_{2}^{\mathrm{min}}$, e.g., it could go to subsolar masses. Future work could update this assumption by imposing a non-zero lower bound on $m_{2}^{\mathrm{min}}$, representing a minimum black hole mass, which may help us understand the classification of GW190814. Additional model flexibility also could be added, for example by relaxing the assumption that $m_{2}^{\mathrm{min}}(m_{\mathrm{min}})=m_{\mathrm{min}}$; Farah:2023swu find that the inferred minimum primary and minimum secondary masses of the BBH population are in agreement within current uncertainties, but did not include GW190814 in their analysis. Our parametric forms for the minimum secondary mass are also not motivated by particular astrophysics or trend in the data. A data-driven description of the primary-mass and mass-ratio distributions may give more insight into the underlying correlations, unrestricted by the choice of parametric form. Furthermore, we specify a power-law distribution in mass ratio, restricting the mass-ratio distribution to be strictly monotonic. However, models with more relaxed assumptions about the shape of the distribution show support for the mass-ratio distribution to peak away from $1$ (Sadiq:2023zee; godfreyCosmicCousinsIdentification2023; Rinaldi:2023bbd; Callister:2023tgi; Rinaldi:2025emt). Other works find that the mass-ratio distribution changes with primary mass, either continuously or for different subpopulations, motivating the exploration of a variable shape of the distribution in addition to a variable lower bound (Li:2022jge; Baibhav:2022qxm; Galaudage:2024meo; Roy:2025ktr). Considering a mass-ratio distribution that more accurately reflects the data would help avoid biases driven by an inaccurate models.
@@
Other observable parameters such as the spin or eccentricity could provide more information about the astrophysical nature of such BBHs. Works looking at the correlations between mass and spin may identify subpopulations of sources from similar evolutionary pathways (e.g., Safarzadeh:2020mlb; Heinzel:2023hlb; Ray:2024hos; Antonini:2024het). If the mass-ratio distribution is jointly correlated with component mass and spin distributions, neglecting to model these correlations could lead to biased inference (Alvarez-Lopez:2025ltt). Jointly investigating mass, mass-ratio, and spin distributions could help us understand how these parameters are correlated, and therefore give more insight into the astrophysical evolution of the population. However, adding additional parameters to our models and investigating the joint correlations between mass, mass ratio, and spin, requires having sufficient data to constrain the model parameters. The increased number of BBH observations coming with the forth LIGO–Virgo–KAGRA observing run will allow us to better measure any correlations between the BBH observables, and explore the astrophysics encoded in these correlations.
@@ acknowledgments
We thank the groups at Northwestern University for hospitality while this work was developed, in particular to Sylvia Biscoveanu, Monica Gallegos-Garcia, and Miguel Martinez for insightful discussions. We thank Alexandra Guerrero for comprehensive review, Daniel Holz for valuable insights, John Veitch for useful technical discussions, and the anonymous referee for helpful comments. SC is supported by Science and Technology Facilities Council (STFC) studentship 2748218. CPLB is supported by STFC grant ST/V005634/1. ZD acknowledges support from the Center for Interdisciplinary Exploration and Research in Astrophysics (CIERA) Board of Visitors Research Professorship. SC thanks the CIERA Board of Visitors and the University of Glasgow MacRobertson Scholarship for support to visit CIERA. This research has made use of data from the Gravitational Wave Open Science Center, a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by the National Science Foundation (NSF) LIGO Laboratory which is a major facility fully funded by the NSF, as well as STFC of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO 600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by NSF Grants PHY-0757058 and PHY-0823459. This document has been assigned LIGO document number LIGO-P2500498.
These results were produced using models found in this code release. Results and data products used in this work are available on Zenodo: https://doi.org/10.5281/zenodo.16895783 (catalog doi:10.5281/zenodo.16895783) (variableqminDR).